OpenStudy (anonymous):

efsvChris ran 4 miles more than Miguel ran, the sum of their distances is 18 miles. How far did Miguel run? The domain of the solution is {04,7,14} a) Define your variable b) Write two equations using the information from the problem c) Use the equations from b) to write ONE equation to help you answer the question given in the problem. d) Substitute each value from the domain into your equation to determine the correct solution (You must substitute ALL values from the domain to get full credit for this part of the problem)

OpenStudy (anonymous):

Something wrong?

OpenStudy (anonymous):

Yeah, x

OpenStudy (anonymous):

... x

OpenStudy (anonymous):

The number of miles Miguel ran, right?

OpenStudy (anonymous):

Okay!

OpenStudy (ikram002p):



OpenStudy (anonymous):

Okay so C= 4 - 18?

OpenStudy (ikram002p):

now let pink line is what Chris ran, make is C... while the red is what miguel did M now C+M=18 agree? (given)

OpenStudy (anonymous):

Yes

OpenStudy (ikram002p):

sure ? *no lying aloud*

OpenStudy (anonymous):

Yes

OpenStudy (ikram002p):

now how ever much migual ran chris ran more by 4 miles which means the difference is 4, how would u write that in equation @SeriousGamer29 ?

OpenStudy (anonymous):

M = 18 - 4

OpenStudy (ikram002p):

:O well we don't know how much chris have ran do we?

OpenStudy (anonymous):

Oh... right...

OpenStudy (ikram002p):